OpenStudy (anonymous):

find the fourth term of the geometric sequence whose first term is 6 and whose common ratio is -3

OpenStudy (amistre64):

-3(-3(-3(-3(6))))

OpenStudy (anonymous):

the pattern is multiplying -3 with each term since we start with 6 we have 1st term=6 2nd term=6*(-3)=-18 3rd term= -18*(-3)=54 4th term= 54*(-3)=-162

OpenStudy (anonymous):

thank u...

OpenStudy (amistre64):

got to many -3 in mine :)

OpenStudy (anonymous):

thank u..